# On Lan-Sheng-Zuo conjecture

### v1

@@ Abstract.
In this paper we study the Lan-Sheng-Zuo conjecture proposed in [LSZ13]. We prove that the conjecture holds for smooth projective curves with genus $g\leq 1$, and construct explicit counter-examples of arbitrary big rank (the first example is $p=2,r=3$) Higgs bundles over any smooth projective curves with genus $g\geq 2$.
Assumption
In this paper, $k$ always denotes an algebraically closed field.
@@ 1. Introduction
Let $C$ be a smooth projective curve over $k$. A classical question in the theory of vector bundles over positive characteristic is whether every semistable vector bundle is strongly semistable or not. D. Gieseker [Gie73] constructed the first example of a semistable rank two vector bundle over $k$ which is not strongly semistable.
In [LSZ13], they introduced the notion of strongly semistability of Higgs bundles, which generalized the notion of strongly semistability of vector bundle naturally, and proved that every rank two semistable Higgs bundle is strongly semistable in ([LSZ13, Theorem 2.6]). Moreover, it was conjectured by Lan-Sheng-Zuo that every semistable Higgs bundle, therefore particularly every semistable bundle, is strongly semistable. The precise statement is given as follows:
@@
When the rank of a vector bundle over $k$ is greater than $p=\textrm{char}(k)$, it is said to be big. It is not much known about the Conjecture 1.1 for semistable Higgs bundles of big rank, and the main results of this paper are summarized as follows:
@@
    If $g(C)\leq 1$, Conjecture 1.1 holds true;
@@
    If $g(C)\geq 2$, then for any big rank $r$, there exist rank $r$ semistable Higgs bundles over $C$ which are not strongly semistable. In particular, Conjecture 1.1 fails for some Higgs bundles of big rank.
@@
Note that for $C=\mathbb{P}^{1}$, Conjecture 1.1 holds trivially, since semistable Higgs bundles on $\mathbb{P}^{1}$ are direct sums of line bundles of the same degree, equipped with the zero Higgs field.
@@
In the first part, we prove that the Simpson filtration can always be chosen such that the associated graded Higgs bundle is polystable (Theorem 3.1.1), and as an application, we prove that the Lan-Sheng-Zuo conjecture is true on the elliptic curves (Proposition 3.2.1).
In the second part, firstly we point out the existence of counter-examples of Lan-Sheng-Zuo conjecture on certain general type varieties with dimension $n\geq 2$ (Corollary 4.2.1), which is an immediate consequence of results in [Zho16]. By the restriction theorem, this implies the existence of counter-examples for Lan-Sheng-Zuo conjecture on generic curves in some linear system. Secondly, we provide an explicit construction of counter-examples of Lan-Sheng-Zuo conjecture for all possible cases on smooth projective curve (Proposition 4.2.1, Remark 4.2.3). In particular, a consequence is that we show that the assumption in Theorem 1.1 is sharp.
In the end of second part, we prove the existence of uniformizing flat bundle in the positive characteristic case (Lemma 4.3.1) and use it to provide a tensor product type of counter-examples of Lan-Sheng-Zuo conjecture, which also implies the category of strongly semistable Higgs bundles fails to be a tensor category.
In the last part, we collect some partial results on Lan-Sheng-Zuo conjecture on Calabi-Yau varieties.
@@ Acknowledgement
The authors would like to express gratitude to Jianping Wang, who motivated the proof of Lemma 4.3.1, and to Xucheng Zhang, for discussions about stack theory. Also, the authors would like to thank Lingguang Li, for his careful reading of the manuscript and for providing invaluable comments that improved the paper.
The second named author is supported by Chinese Academy of Sciences Project for Young Scientists in Basic Research (Grant No. YSBR-032).
@@ 2. $(2)$
The Ogus and Vologodsky established the non-abelian Hodge theory in positive characteristic case, and a simple version can be stated as follows:
@@
The inverse Cartier of a vector bundle $\mathcal{E}$ is the Frobenius pullback $F^{*}\mathcal{E}$. In the Higgs field non-trivial case, the inverse Cartier transform can be explained as an exponential twisting of Frobenius map by Higgs field in [LSZ15].
@@ 3.1.1. Motivation and first examples
In [Lan14], Langer claimed that every Simpson filtration of a flat bundle $(\mathcal{V},\nabla)$ can be refined such that the associated graded Higgs bundle is polystable. However, we point out that not every Simpson filtration can be refined to be a gr-polystable filtration.
@@
Motivated by above example, instead of considering refinement for any Simpson filtration, we should prove that there always exists a Simpson filtration such that the associated graded Higgs bundle is polystable, which is called a gr-polystable filtration for convenience.
@@
Let $E$ be an elliptic curve and $(\mathcal{V},\nabla)$ be $\nabla$-semistable flat bundle on $E$. Recall that every $\nabla$-semistable flat bundle on elliptic curve is semistable as a vector bundle, so $\mathcal{V}$ is a semistable vector bundle. Thus an obvious Simpson filtration of $(\mathcal{V},\nabla)$ is the trivial filtration. In this case, the gr-polystable filtration is given by the Jordan-Hölder filtration of $\mathcal{V}$.
3.1.2. Existence of gr-polystable filtration
In this section, we prove that every $\nabla$-semistable flat bundle over a smooth projective variety $X$ over $k$ has a Simpson filtration such that the associated graded Higgs bundle is polystable by an argument based on algebraic stacks.
Let $\Theta:=[\operatorname{\mathbb{A}}^{1}/\operatorname{\mathbb{G}}_{m}]$ be the quotient stack defined by the standard contracting action of the multiplicative group on the affine line $\operatorname{\mathbb{A}}^{1}$ and $\mathfrak{X}$ be an algebraic stack. A morphism $f\colon\Theta_{k}\to\mathfrak{X}$ is the same as a $\operatorname{\mathbb{G}}_{m}$-equivariant bundle on $\operatorname{\mathbb{A}}^{1}$, and by Rees construction it can be thought as an object of $x_{1}\in\mathfrak{X}(k)$, which is the object $f(1)$, together with a filtration of $x_{1}$ and as $f(0)=x_{0}$ as the associated graded object. The following lemma is the key technique we will use.
Lemma 3.1.1 ([AHLH23, Lemma 3.24]).
Let $\mathfrak{X}$ be a locally reductive algebraic stack locally of finite type over $k$. Then any specialization $x\rightsquigarrow x_{0}$ of $k$-points is realized by a morphism $\Theta_{k}\to\mathfrak{X}$, where $x_{0}$ is a closed point.
Let $\mathfrak{R}_{\text{Hod}}$ be the algebraic stack of the $\nabla$-semistable flat $\lambda$-connections on a smooth projective variety $X$ over $k$. By using techniques in the proof of [WZ24, Proposition 3.28], we have $\mathfrak{R}_{\text{Hod}}$ is a locally reductive algebraic stack locally of finite type over $k$ since it admits a moduli space. There is a natural map $\pi\colon\mathfrak{X}\to\operatorname{\mathbb{A}}^{1}$ with $\pi^{-1}(0)=\mathfrak{R}_{\text{Hig}}$ is the stack of semistable Higgs bundles and $\pi^{-1}(1)=\mathfrak{R}_{\text{dR}}$ is the stack of semistable flat bundles.
Theorem 3.1.1.
Let $X$ be a smooth projective variety and $\omega$ be an ample divisor. Let $(\mathcal{E},\theta)$ be a $\omega$-semistable Higgs bundle over $X$ which is nilpotent of exponent $\leq p-1$ and $(\mathcal{V},\nabla):=C^{-1}(\mathcal{E},\theta)$. Then there exists a Simpson filtration on $(\mathcal{V},\nabla)$ such that the associated graded Higgs bundle is polystable.
Proof.
Firstly, there already exists a filtration $\mathcal{N}_{S}^{\bullet}$ for a $\nabla$-semistable flat bundle $(\mathcal{V},\nabla)$ such that the associated graded Higgs bundle $(\mathcal{E},\theta)$ is semistable. By Rees construction, it gives a $\operatorname{\mathbb{G}}_{m}$-equivariant family in $\mathfrak{R}_{\text{Hod}}$ over $\operatorname{\mathbb{A}}^{1}$, such that object over $1\in\operatorname{\mathbb{A}}^{1}$ is $(\mathcal{V},\nabla)$ and the object over $0\in\operatorname{\mathbb{A}}^{1}$ is $(\mathcal{E},\theta)$. In particular, $(\mathcal{V},\nabla)$ specializes to $(\mathcal{E},\theta)$ in $\mathfrak{R}_{\text{Hod}}$.
On the other hand, there is a unique polystable representative $(\mathcal{E}^{\prime},\theta^{\prime})$ in the $S$-equivalent class of $(\mathcal{E},\theta)$, which is the associated graded bundle of $(\mathcal{E},\theta)$ with respect to the Jordan-Hölder filtration. Then by Rees construction it gives a $\operatorname{\mathbb{G}}_{m}$-equivariant family in $\mathfrak{R}_{\text{Hig}}$. In particular, $(\mathcal{E},\theta)$ specializes to $(\mathcal{E}^{\prime},\theta^{\prime})$ in $\mathfrak{R}_{\text{Hig}}$
Combining these two specializations together we get a specialization $(\mathcal{V},\nabla)\leadsto(\mathcal{E}^{\prime},\theta^{\prime})$ in $\mathfrak{R}_{\text{Hod}}$. By Lemma 3.1.1, this specialization can be realized by a map $f\colon\Theta_{k}\to\mathfrak{R}_{\text{Hod}}$ with $f(1)=(\mathcal{V},\nabla)$ and $f(0)=(\mathcal{E}^{\prime},\theta^{\prime})$. In other words, this shows that there exists a filtration on $(\mathcal{V},\nabla)$ such that the associated graded Higgs bundle is polystable. ∎
@@ 3.2. Lan-Sheng-Zuo conjecture on elliptic curves
In this section, we provide two different proofs of Lan-Sheng-Zuo conjecture on elliptic curve $E$. One of the proofs is based on the existence of gr-polystable filtration and Lemma 3.2.1, and another proof is based on the fact that every semistable bundle on elliptic curve is strongly semistable.
@@
Let $(\mathcal{V},\nabla)=C^{-1}(\mathcal{E},\theta)$. Since $\mathcal{V}$ is semistable as a vector bundle, we may choose trivial filtration on $(\mathcal{V},\nabla)$ as its Simpson filtration, then the associated graded Higgs bundle is $(\mathcal{V},0)$. Since every semistable bundle on elliptic curve is strongly semistable, $(\mathcal{E},\theta)$ initiates a semistable Higgs de Rham flow as follows
@@ 4.3.2. The construction of tensor product type counter-example
5. Lan-Sheng-Zuo conjecture on Calabi-Yau varieties
The technique used in the second proof of Proposition 3.2.1 can be generalized to higher dimensional cases, according to the following several results.
Theorem 5.1 ([Lan15, Theorem 0.1]).
@@ 5. Lan-Sheng-Zuo conjecture on Calabi-Yau varieties
Theorem 5.2 ([Lan15, Theorem 4.5]).
Let $X$ be a K3 surface over an algebraically closed field. Then the tangent bundle of $X$ is strongly semistable with respect to some non-trivial nef polarization.
Proposition 5.1.
Let $X$ be a smooth $n$-dimensional projective variety over $k$ and $\omega$ be a polarization on $X$. Assume that $K_{X}$ is numerically trivial and $\omega$-strongly semistable. Then any $\omega$-semistable Higgs bundle $(\mathcal{E},\theta)$ or $\nabla$-semistable flat bundle $(\mathcal{V},\nabla)$ on $X$ is $\omega$-semistable as a vector bundle.
Proof.
For convenience here we only give the proof for the Higgs bundle case. Suppose $\mathcal{E}$ is unstable and its Harder-Narasimhan filtration is given by
$0=\mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\dots\subset\mathcal{E}_{s-1}\subset\mathcal{E}_{s}=\mathcal{E},$
where $\mathcal{G}_{i}=\mathcal{E}_{i}/\mathcal{E}_{i-1}$ is $\omega$-semistable bundle for each $i$ with strictly decreasing slopes. Since $\mathcal{E}$ is $\omega$-semistable if $(F^{*})^{N}\mathcal{E}$ is $\omega$-semistable, so up to finite times Frobenius pullback, we may assume $\mathcal{G}_{i}$ is strongly $\omega$-semistable for each $i$.
Suppose the maximally destabilizer $\mathcal{E}_{1}$ is not $\theta$-invariant and take $\ell>1$ such that $\theta\left(\mathcal{E}_{1}\right)\subseteq\mathcal{E}_{\ell}$ but $\theta\left(\mathcal{E}_{1}\right)\nsubseteq\mathcal{E}_{\ell-1}$. Then $\theta$ induces a non-zero morphism $\mathcal{G}_{1}\to\mathcal{G}_{\ell}\otimes\Omega_{X}^{1}$. But if $\ell\neq 1$, there is no non-zero morphism between $\mathcal{G}_{1}$ and $\mathcal{G}_{\ell}\otimes\Omega_{X}^{1}$ since $K_{X}$ is numerically trivial and both $\mathcal{G}_{1}$ and $\mathcal{G}_{\ell}\otimes\Omega_{X}^{1}$ are semistable. Then $\mathcal{E}_{1}$ is $\theta$-invariant and $(\mathcal{E},\nabla)$ is not a semistable Higgs bundle, a contradiction. ∎

### v2

@@ Abstract.
In this paper we study the Lan-Sheng-Zuo conjecture proposed in [LSZ13]. We prove that the conjecture holds for smooth projective curves with genus $g\leq 1$, and construct explicit counter-examples of arbitrary big rank (the first example is $p=2,r=3$) Higgs bundles over any smooth projective curve with genus $g\geq 2$.
@@ 1. Introduction
In this paper, unless otherwise specified, $k$ always denotes an algebraically closed field with $\text{char}(k)=p>0$.
Let $C$ be a smooth projective curve over $k$. A classical question in theory of vector bundles over positive characteristic is whether every semistable vector bundle is strongly semistable or not. D. Gieseker [Gie73] constructed the first example of a semistable rank two vector bundle on a smooth projective curve over $k$ which is not strongly semistable.
In [LSZ13], Lan, Sheng and Zuo introduced the notion of strongly semistability of Higgs bundles, which generalized the notion of strongly semistability of vector bundle naturally, and proved that every rank two semistable Higgs bundle is strongly semistable ([LSZ13, Theorem 2.6]). Moreover, it was conjectured by Lan-Sheng-Zuo that every semistable Higgs bundle, therefore particularly every semistable bundle, is strongly semistable. The precise statement is given as follows:
@@
When the rank of a vector bundle over $k$ is greater than $p$, it is said to be big. It is not much known about the Conjecture 1.1 for semistable Higgs bundles of big rank, and the main results of this paper are summarized as follows:
@@
    If $g(C)\leq 1$, then Conjecture 1.1 holds true;
@@
    If $g(C)\geq 2$, then for any big rank $r$, there exists a rank $r$ semistable Higgs bundle over $C$ which are not strongly semistable. In particular, Conjecture 1.1 fails for some Higgs bundles of big rank.
@@
Note that for $C=\mathbb{P}^{1}$, Conjecture 1.1 holds trivially, as semistable Higgs bundles on $\mathbb{P}^{1}$ are direct sums of line bundles of the same degree, equipped with the zero Higgs field.
@@
In the first part, we prove that the Lan-Sheng-Zuo conjecture holds on the elliptic curves (Proposition 3.2.1).
In the second part, firstly we show the existence of counter-examples of the Lan-Sheng-Zuo conjecture on certain varieties of general type with dimension $n\geq 2$ (Corollary 4.2.1), which is an immediate consequence of results in [Zho16]. By restriction theorem, this implies the existence of counter-examples for the Lan-Sheng-Zuo conjecture on generic curves in some linear system. After that, we provide an explicit construction of counter-examples of Lan-Sheng-Zuo conjecture for all possible cases on smooth projective curve (Proposition 4.2.1, Remark 4.2.3). In particular, an immediate consequence is that the assumption in Theorem 1.1 is sharp.
In the end of second part, we prove the existence of uniformizing flat bundle in the positive characteristic case (Lemma 4.3.1) and use it to provide a tensor product type of counter-examples of the Lan-Sheng-Zuo conjecture, which also provides an example to show the category of strongly semistable Higgs bundles is not a tensor category.
In the last part, we collect some partial results of the Lan-Sheng-Zuo conjecture on higher dimensional varieties.
@@ Acknowledgement
The authors would like to express gratitude to Jianping Wang, who motivated the proof of Lemma 4.3.1 and Proposition 5.1, and to Xucheng Zhang, for discussions about stack. Also, the authors would like to thank Lingguang Li, for his careful reading of the manuscript and for providing invaluable comments that improved the paper.
The second named author is supported by the Chinese Academy of Sciences Project for Young Scientists in Basic Research (Grant No. YSBR-032).
@@ 2. $(2)$
A. Ogus and V. Vologodsky established the non-abelian Hodge theory in positive characteristic case, and a simple version can be stated as follows:
@@
The inverse Cartier of a vector bundle $\mathcal{E}$ is the Frobenius pullback $F^{*}\mathcal{E}$. In the non-trivial Higgs field case, the inverse Cartier transform can be explained as an exponential twisting of Frobenius map by Higgs field in [LSZ15].
@@ 3.1.1. Motivation and first examples
In [Lan14], A. Langer claimed that every Simpson filtration of a flat bundle $(\mathcal{V},\nabla)$ can be refined such that the associated graded Higgs bundle is polystable. However, we point out that not every Simpson filtration can be refined to be a gr-polystable filtration.
@@
Instead of considering refinement for an arbitrary Simpson filtration, we should prove that there always exists a Simpson filtration such that the associated graded Higgs bundle is polystable, which is called a gr-polystable filtration for convenience. For smooth projective curve of genus zero and one, it is easy to figure out which one is gr-polystable filtration.
@@
Let $E$ be an elliptic curve and $(\mathcal{V},\nabla)$ be $\nabla$-semistable flat bundle on $E$. Since every $\nabla$-semistable flat bundle on an elliptic curve is semistable as a vector bundle, so $\mathcal{V}$ is a semistable vector bundle. Thus an obvious Simpson filtration of $(\mathcal{V},\nabla)$ is the trivial filtration, and the gr-polystable filtration is given by the Jordan-Hölder filtration of $\mathcal{V}$.
@@ 3.2. Lan-Sheng-Zuo conjecture on elliptic curves
In this section, we provide two different proofs of Lan-Sheng-Zuo conjecture on an elliptic curve $E$. One of the proofs is based on the existence of gr-polystable filtration and Lemma 3.2.1, and another proof is based on the fact that every semistable bundle on an elliptic curve is strongly semistable.
@@
Let $(\mathcal{V},\nabla)=C^{-1}(\mathcal{E},\theta)$. Since $\mathcal{V}$ is semistable as a vector bundle, we may choose trivial filtration on $(\mathcal{V},\nabla)$ as its Simpson filtration, then the associated graded Higgs bundle is $(\mathcal{V},0)$. Since every semistable bundle on an elliptic curve is strongly semistable, $(\mathcal{E},\theta)$ initiates a semistable Higgs de Rham flow as follows
@@ 4.3.2. The construction of tensor product type counter-example
5. Lan-Sheng-Zuo conjecture on higher dimensional varieties
In this section we collect some partial results on Lan-Sheng-Zuo conjecture on higher dimensional varieties. Let $X$ be a smooth $n$-dimensional projective variety over $k$ and $\omega$ be an ample polarization on $X$.
If $\mu_{\omega,\max}(\Omega_{X}^{1})\leq 0$, then by Theorem 4.1.1, every $\omega$-semistable vector bundle is strongly semistable. Thus a sufficient method to prove Lan-Sheng-Zuo conjecture on $X$ is to prove every $\nabla$-semistable flat bundle $(\mathcal{V},\nabla)$ is semistable as a vector bundle.
One method to prove every $\nabla$-semistable flat bundle $(\mathcal{V},\nabla)$ is semistable as bundle is to use following estimates:
Lemma 5.1 ([Lan15a, Lemma 5]).
Let $(\mathcal{V},\nabla)$ be a $\nabla$-semistable flat bundle of rank $r$. Then
$L_{\max}(\mathcal{E})-L_{\min}(\mathcal{E})\leq(r-1)\gamma,$
where
$L_{\omega,\max}(\mathcal{E})$ $=\lim_{t\to\infty}\frac{\mu_{\omega,\max}((F^{*})^{k}\mathcal{E})}{p^{k}}$
$L_{\omega,\min}(\mathcal{E})$ $=\lim_{t\to\infty}\frac{\mu_{\omega,\min}((F^{*})^{k}\mathcal{E})}{p^{k}}.$
and $\gamma=(r-1)\max\{L_{\max}(\Omega_{X}^{1}),0\}$.
Lemma 5.2 ([Lan04, Corollary 6.2]).
If $p\geq n=\dim X$ and $L_{\max}(\Omega_{X}^{1})\geq 0$, then
$L_{\max}(\Omega_{X}^{1})\leq\frac{p}{p+1-n}\mu_{\max}(\Omega_{X}^{1}).$
Proposition 5.1.
Let $X$ be a smooth $n$-dimensional projective variety over $k$ and $\omega$ be an ample polarization on $X$ such that $\mu_{\omega,\max}(\Omega_{X}^{1})\leq 0$.
  1. $(1)$
    If $\Omega_{X}^{1}$ is strongly $\omega$-semistable, then Lan-Sheng-Zuo holds on $X$.
  2. $(2)$
    If $p\geq n$, then Lan-Sheng-Zuo conjecture holds on $X$.
Proof.
Firstly, we prove under assumptions $(1)$ and $(2)$, we must have $L_{\omega,\max}(\Omega_{X}^{1})\leq 0$. Indeed, for the case (1), if $\Omega_{X}^{1}$ is strongly $\omega$-semistable, then $L_{\omega}(\mathcal{E})=\mu_{\omega}(\mathcal{E})\leq 0$. For the case (2), if $p\geq n$ and $L_{\omega,\max}(\Omega_{X}^{1})>0$, then by Lemma 5.2 we have
$0<L_{\max}(\Omega_{X}^{1})\leq\frac{p}{p+1-n}\mu_{\max}(\Omega_{X}^{1})\leq 0,$
a contradiction.
If $L_{\omega,\max}(\Omega_{X}^{1})\leq 0$, then for any $\nabla$-semistable flat bundle $(\mathcal{V},\nabla)$, by Lemma 5.1 we have $\mu_{\omega,\max}(\mathcal{V})-\mu_{\omega,\min}(\mathcal{V})=0$ since we always have
$0\leq\mu_{\omega,\max}(\mathcal{V})-\mu_{\omega,\min}(\mathcal{V})\leq L_{\omega,\max}(\mathcal{V})-L_{\omega,\min}(\mathcal{V}).$
Then $\mathcal{V}$ is semistable and thus by Theorem 4.1.1 it is strongly semistable. As a consequence, Lan-Sheng-Zuo conjecture holds on $X$ since we can always choose trivial filtration to run the Higgs-de Rham flow. ∎
For the case (1) in Proposition 5.1, it can be also proved by the same technique used in the second proof of Proposition 3.2.1.
Another proof of (1) in Proposition 5.1.
Let $(\mathcal{V},\nabla)$ be a $\nabla$-semistable flat bundle on $X$. Suppose $\mathcal{V}$ is unstable and its Harder-Narasimhan filtration is given by
$0=\mathcal{V}_{0}\subset\mathcal{V}_{1}\subset\dots\subset\mathcal{V}_{s-1}\subset\mathcal{V}_{s}=\mathcal{V},$
where $\mathcal{G}_{i}=\mathcal{V}_{i}/\mathcal{V}_{i-1}$ is $\omega$-semistable bundle for each $i$ with strictly decreasing slopes. Since $\mathcal{V}$ is $\omega$-semistable if $(F^{*})^{N}\mathcal{V}$ is $\omega$-semistable, so up to finite times Frobenius pullback, we may assume $\mathcal{G}_{i}$ is strongly $\omega$-semistable for each $i$.
Since $(\mathcal{V},\nabla)$ is $\nabla$-semistable, we must have the maximally destabilizer $\mathcal{V}_{1}$ is not $\nabla$-invariant. Choose $\ell>1$ such that $\nabla\left(\mathcal{V}_{1}\right)\subseteq\mathcal{V}_{\ell}$ but $\nabla\left(\mathcal{V}_{1}\right)\nsubseteq\mathcal{V}_{\ell-1}$. Then $\nabla$ induces a non-zero morphism $\mathcal{G}_{1}\to\mathcal{G}_{\ell}\otimes\Omega_{X}^{1}$. But if $\ell\neq 1$, there is no non-zero morphism between $\mathcal{G}_{1}$ and $\mathcal{G}_{\ell}\otimes\Omega_{X}^{1}$ since $\mu_{H}(\Omega_{X}^{1})\leq 0$ and both $\mathcal{G}_{1}$ and $\mathcal{G}_{\ell}\otimes\Omega_{X}^{1}$ are semistable, as tensor product of strongly semistable vector bundle is again strongly semistable. This shows $\mathcal{V}_{1}$ is $\nabla$-invariant, a contradiction. ∎
Theorem 5.1 ([Lan15b, Theorem 0.1]).
@@ 2. $(2)$
Theorem 5.2 ([Lan15b, Theorem 4.5]).
Let $X$ be a K3 surface over $k$. Then the tangent bundle of $X$ is strongly semistable with respect to some non-trivial nef polarization.